OpenStudy (m_m13):

will medal what is the best way to classify each question? A) Identity B) Contradiction C) Neither

OpenStudy (m_m13):

1. 5X - 20 = 2(3X - 10) 2. 4X - 16 - 2X = 2(X-8)

OpenStudy (m_m13):

3. 6X - 15 = 3(2x - 6) 4. 5X -3(X -4) = 2(X - 6)

OpenStudy (anonymous):

Do you know what identits and contradictions are in mathematic terms?

OpenStudy (kweenlexis):

0.0 lets see

OpenStudy (m_m13):

help?

OpenStudy (anonymous):

An identity is when the equation is true, such as 3 = 1 + 2. A contradiction is when the equation is not true, such as 3 = 4 x 1. If you simplify the equations given, you should be able to solve whether or not they are identities or contradictions

OpenStudy (m_m13):

@kweenlexis help?

OpenStudy (m_m13):

help simplify

OpenStudy (anonymous):

1. 5X - 20 = 2(3X - 10) If we simplify the second half, we will be able to solve the problem. The only step we need to do with this one is distribute, so we just multiply that which is in the parenthesis by the number outside, in this case 2. 1. 5x - 20 = 6x - 20 is the simplified version for the first problem

OpenStudy (m_m13):

ok...

OpenStudy (anonymous):

As we can see, 5x - 20 does not equal 6x - 20, so that would be a contradiction

OpenStudy (m_m13):

ok

OpenStudy (anonymous):

2. 4X - 16 - 2X = 2(X-8) 4x and -2x are like terms, so we can combine them together to create 2x. 2. 2x - 16 = 2(x - 8) when we use the distributive property, we can simplify the second half to 2. 2x - 16 = 2x - 16 So is this an identity or a contradiction

OpenStudy (kweenlexis):

oh my bad i was doing some thing :(

OpenStudy (m_m13):

its ok

OpenStudy (m_m13):

so 2. is a contradiction

OpenStudy (m_m13):

cause the sides are the same yet they're opposite

OpenStudy (m_m13):

wait is 2. identity

OpenStudy (anonymous):

Yes, 2. is an identity. Are you starting to get the hang of it?

OpenStudy (m_m13):

ya, I just saw the question wrong

OpenStudy (m_m13):

can you help me with the others?

OpenStudy (anonymous):

sure

OpenStudy (m_m13):

thanks i'll give you a medal after

OpenStudy (m_m13):

hi

OpenStudy (m_m13):

hello?

OpenStudy (anonymous):

3. 6X - 15 = 3(2x - 6) this is just a one step distributive, so the second half will be 6x - 15 = 6x - 18 Identity or contradiction? 4. 5X -3(X -4) = 2(X - 6) This one is more complicated, as there are parentheses on both sides. We just need to tackle them one at a time. The first half will come out as 5x - 3x - 12, and then we can combine the like terms (5x and 3x) so that the first half will be 2x - 12 The second half is just distribution, so 2x - 12. The equation ends up simplifying to 2x - 12 = 2x - 12 Identity or contradiction? Hope I helped!

OpenStudy (m_m13):

3. is Identity

OpenStudy (m_m13):

sorry I meant for 4

OpenStudy (m_m13):

4. is identity and 3. is...

OpenStudy (m_m13):

3. is contradiction or neither?

OpenStudy (anonymous):

3. is a contradiction. Good job!

OpenStudy (m_m13):

thank you